Unless stated otherwise, interest is compounded annually, and payments occur at the end of the period. Face value for bonds is $1000.

  1. Brake Plus has a stock price of $30 per share with 12 million shares outstanding. There is 120 million in debt with a yield on debt of 4.6%. The equity beta for Brake is 1.20. The risk-free rate is 2.5% and the market risk premium is 6%. Carry all work to two decimal points (so XX.XX%)
    1. Use the equity beta in the CAPM to get the cost of equity. Use the cost of equity and the yield on the debt to calculate the unlevered cost of capital.
    2. Assume Brake’s debt has a beta of zero. Calculate the unlevered beta. Use unlevered beta in the CAPM to get the unlevered cost of capital.
    3. Why are the two answers different?
Answer #1

formula:

Capital Asset pricing model:

As per CAPM model:
Re= Rf+(Rm-Rf)B

Re= required rate of return.
Rf= Risk-free rate.
Rm = Return on market.
Rm-Rf =Market Risk Premium.
B = Beta, systematic risk.

Answers:
1. Levered Cost of capital = 9.7%
2. Unlevered Beta = 0.9
Unlevered cost of capital is 7.9%
3. The difference arises due to the debt component. Which can be clearly seen in the asset Beta formula.

a) levend Beta LRe= RA+ (Rm-Rp) ße 1 = 2.5+ 6x1.20 = 9.7% bez leverd Betal Equity beta- b) 13a = ße [ite ci-t) D-debt Ez Equi